OpenStudy (anonymous):

If the point P(-3/5,y) lies on the unit circle and P is in the third quadrant, what does y equal? If necessary, use the slash mark ( / ) for a fraction bar.

OpenStudy (ranga):

The equation for unit circle is x^2 + y^2 = 1 The point P has x = -3/5. Plug in x = -3/5 and solve for y. Remember that P lies in the third quadrant which means both x and y values on the circle will be negative. So your y answer must be negative.

OpenStudy (goformit100):

"A warm Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

So -4/5

OpenStudy (ranga):

Yes!